# FLUID: A Fine-Grained Lightweight Urban Signalized-Intersection Dataset of Dense Conflict Trajectories

### v1

@@ Acknowledgements
We are grateful to the Joint Research and Development Laboratory of Smart Policing in Xuancheng Public Security (a joint initiative between Xuancheng Public Security Bureau and Sun Yat-sen University) for their essential contribution to this work, specifically in facilitating the selection of the investigation’s spatiotemporal scope and data acquisition. We also thank several anonymous undergraduate students from the School of Intelligent System Engineering at SYSU for assisting the authors in completing data integration. The research was also supported by the National Key Research and Development Program of China (No. 2023YFB4301900), the Joint Funds of the National Natural Science Foundation of China (No. U21B2090), the General Program Sponsored by Shenzhen Natural Science Foundation in Basic Research Fund (No.JCYJ20240813151445059), and the Science and Technology Planning Project of Guangdong Province (No. 2023B1212060029).
@@
Y.C., Z.W. and G.Z. conceived and designed the study. Y.C., Z.W. and L.X. collected the data. Y.C. and G.Z. performed the experiments and analyzed the results, which were validated by L.X. and H.T. Y.C. wrote the original manuscript, which was reviewed and edited by Z.W. and X.W. Z.H. supervised the project and provided guidance. H.Z. and Z.H. were responsible for project administration and acquired funding. All authors reviewed the manuscript.

### v2

@@ Acknowledgements
We are grateful to the Joint Research and Development Laboratory of Smart Policing in Xuancheng Public Security (a joint initiative between Xuancheng Public Security Bureau and Sun Yat-sen University) for their essential contribution to this work, specifically in facilitating the selection of the investigation’s spatiotemporal scope and data acquisition. We thank Junfeng Li and Kaiyuan Yang for their assistance in vector map processing, as well as the anonymous undergraduate students from the School of Intelligent System Engineering at SYSU for their help with data integration. The research was also supported by the National Key Research and Development Program of China (No. 2023YFB4301900), the National Natural Science Foundation of China (No. U21B2090), the Shenzhen Science and Technology Program (No.JCYJ20240813151445059), and the Science and Technology Planning Project of Guangdong Province (No. 2023B1212060029).
@@
Conceptualization and Methodology: Y.C., Z.W., G.Z., H.Z.; Investigation: Y.C., Z.W., L.X.; Data Curation and Formal Analysis: Y.C., G.Z.; Validation and Visualization: Y.C., L.X., H.T., Z.W.; Writing – Original Draft Preparation: Y.C.; Writing – Review & Editing: Z.W, X.W., Z.H.; Project Administration and Funding Acquisition: H.Z., Z.H. All authors have reviewed and agreed to the manuscript.